Originally scheduled as science run but because of remaining summer shutdown DX AdSec changes and SX Soul still engineering necessary.

Log

When What Log Notes
20190113 DX checkouts DayTime20190113 Closed for snow
20180114 DX checkouts, observation DayTime20190114 Early night cleared up, smooth operation
20190115 DX calibration DayTime20190115 DX calibration KL_v26 up to 150 modes all night.
20190116 Observation DayTime20190116 Observed with KL_v26, found LM1-DX heating not working.
20190117 Observation DayTime20190117 Closed for clouds
20190118 Observation DayTime20190118 Cloudes and frost on the telescope until about UT 3:00, open for the rest of the night. Heater still not working.
20190119 Observation DayTime20190119 Open but fighting MOS error, not working heater, ASM rips, etc.
-- WolfgangGaessler - 16 Jan 2019
Topic revision: r5 - 20 Jan 2019, LorenzoBusoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback